# Queryable Prototype Multiple Instance Learning with Vision-Language Models for Incremental Whole Slide Image Classification

### v1

@@ Abstract
Whole Slide Image (WSI) classification has very significant applications in clinical pathology, e.g., tumor identification and cancer diagnosis. Currently, most research attention is focused on Multiple Instance Learning (MIL) using static WSI datasets. One of the most obvious weaknesses of these methods is that they cannot efficiently preserve and utilize previously learned knowledge. With any new data arriving, classification models are required to be re-trained on both previous and current new data. To overcome this shortcoming and break through traditional vision modality, this paper proposes the first Vision-Language-based framework with Queryable Prototype Multiple Instance Learning (QPMIL-VL) specially designed for incremental WSI classification. This framework mainly consists of two information processing branches: one is for generating bag-level features by prototype-guided aggregation of instance features, while the other is for enhancing class features through a combination of class ensemble, tunable vector and class similarity loss. The experiments on four public WSI datasets demonstrate that our QPMIL-VL framework is effective for incremental WSI classification and often significantly outperforms other compared methods, achieving state-of-the-art (SOTA) performance.
@@ Introduction
Histopathology Whole Slide Image (WSI) is crucial for the diagnosis and treatment of tumor diseases (Bera et al. 2019; Song et al. 2023). To model gigapixel WSIs (e.g., $000\times 90,$ 000 × 90 , 000 pixels) for training clinical-grade models, many methods based on Multiple Instance Learning (MIL) have been extensively studied in the field of computational pathology (CPATH) (Campanella et al. 2019; Lu et al. 2021; Lin et al. 2023, 2024; Liu et al. 2024c). A shared characteristic of these methods is that they are specially designed for capturing a static data distribution (e.g., a given WSI dataset with two lung cancer subtypes), still following conventional machine learning paradigm. However, in the real-world, the distribution of WSI data could be dynamic due to the arrival of new datasets or the discovery of emerging cancer types (Van der Laak, Litjens, and Ciompi 2021; Derakhshani et al. 2022). As a result, a well-trained WSI classification model that is suitable for previous data often cannot adapt to new data. Therefore, the model must be re-trained on the entire dataset (including both previous and new data), leading to a significant increase in training costs. This presents the necessities of new methods well-tailored to dynamic data distributions.
@@
Recent works, e.g., PLIP (Huang et al. 2023b), Prov-GigaPath (Xu et al. 2024), and CONCH (Lu et al. 2024), have deeply demonstrated the significant potential of Vision-Language Models (VLMs) in CPATH. Breaking through traditional vision modality, this kind of new models could effectively exploit pathology features from both vision and language modalities. All of these motivate us to leverage foundational pathology VLMs to study incremental learning for WSI classification.
@@
III) Extensive comparison and ablation experiments are performed on four public WSI datasets. The results demonstrate the superiority of our QPMIL-VL over the existing methods.
@@
Prototype Pool. Specifically, the prototype pool consists of $M$ (key, prompt) prototype pairs $\{(\boldsymbol{k}_{i},\boldsymbol{P}_{i})\}_{i=1}^{M},\boldsymbol{k}_{i}\in\mathbb{R}^{D_{f}},\boldsymbol{P}_{i}\in\mathbb{R}^{L_{\boldsymbol{P}}\times D_{e}}$, where $D_{f}$ is the output feature dimension of encoder $\boldsymbol{E}_{\text{img}}$ and $\boldsymbol{E}_{\text{txt}}$, $D_{e}$ represents the embedding dimension of $\boldsymbol{E}_{\text{txt}}$ and $L_{\boldsymbol{P}}$ indicates the total number of learnable vectors in $\boldsymbol{P}_{i}$. In a prototype pair, the key works as a prototype identifier, used by WSIs for querying, while the prompt acts as a prototype descriptor, utilized to describe the specific visual feature of instance prototype within WSIs.
@@
where $Top\text{-}N^{min}$ denotes the operation of choosing the top-$N$ prototype keys with minimal $\boldsymbol{z}$-key distances. However, in prototype matching on different WSI datasets by Eq. 1, we observe that almost all samples tend to choose the same set of $N$ prototypes from pool. This could usually lead to catastrophic forgetting in incremental learning. To effectively avoid this tendency, we define a penalty factor $p_{i}^{t_{c}}$ to optimize prototype choosing:
@@
FineTune (Lower) 0.308 $\pm$ 0.045 0.336 $\pm$ 0.045 0.841 $\pm$ 0.054 -0.841 $\pm$ 0.054 0.844 $\pm$ 0.038
Regularization-based EWC (Kirkpatrick et al. 2017) 0.309 $\pm$ 0.051 0.338 $\pm$ 0.055 0.836 $\pm$ 0.064 -0.836 $\pm$ 0.064 0.502 0.840 $\pm$ 0.035
LwF (Li and Hoiem 2017) 0.378 $\pm$ 0.081 0.412 $\pm$ 0.083 0.731 $\pm$ 0.116 -0.731 $\pm$ 0.116 0.916 $\pm$ 0.031
Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.436 $\pm$ 0.058 0.477 $\pm$ 0.070 0.670 $\pm$ 0.072 -0.670 $\pm$ 0.072 0.502 0.879 $\pm$ 0.045
ER-ACE/30 (Caccia et al. 2021) 0.666 $\pm$ 0.049 0.732 $\pm$ 0.051 0.075 $\pm$ 0.057 -0.070 $\pm$ 0.059 0.917 $\pm$ 0.023
DER++/30 (Buzzega et al. 2020) 0.749 $\pm$ 0.055 0.823 $\pm$ 0.059 0.219 $\pm$ 0.059 -0.214 $\pm$ 0.062 0.904 $\pm$ 0.029
ER/30 (Chaudhry et al. 2019) 0.790 $\pm$ 0.040 0.870 $\pm$ 0.049 0.186 $\pm$ 0.044 -0.186 $\pm$ 0.044 0.894 $\pm$ 0.036
@@
ER/100 (Chaudhry et al. 2019) 0.827 $\pm$ 0.028 0.910 $\pm$ 0.034 0.143 $\pm$ 0.034 -0.142 $\pm$ 0.032 0.502 0.918 $\pm$ 0.028
Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.616 $\pm$ 0.056 0.677 $\pm$ 0.071 0.285 $\pm$ 0.059 -0.285 $\pm$ 0.059 0.111 0.844 $\pm$ 0.026
MI-Zero (Lu et al. 2023) 0.839 $\pm$ 0.034 0.927 $\pm$ 0.044 - - - 0.909 $\pm$ 0.019
PQMIL-VL (ours) 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014 0.365 0.930 $\pm$ 0.018
Table 1: Main results in class-incremental learning scenario on forward-order training. The best performances are highlighted as bold. “Params” represents the number of learnable parameters. Gray numbers are directly cited from published paper. “/30” and “/100” mean buffer $30$ WSIs and buffer $100$ WSIs, respectively. We present the Masked ACC reflecting task-incremental learning scenario for reference.
@@
To optimize our QPMIL-VL model in training, based on the three losses above, total loss $\mathcal{L}_{O}$ is computed by
$\mathcal{L}_{O}=\mathcal{L}_{C}+\lambda\cdot\mathcal{L}_{M}+\beta\cdot\mathcal{L}_{S},$ (10)
where balance factors $\lambda$ and $\beta$ are set to constant $0.5$. Therefore, our optimization objective is to minimize $\mathcal{L}_{O}$.
@@
Following the setup of previous representative work (Van de Ven and Tolias 2019), we mainly evaluate QPMIL-VL in the class-incremental learning scenario and additionally report the results in the task-incremental scenario.
@@
Compared Methods. It is often necessary to compare with the upper and lower bounds of classification methods. We employ the classic ABMIL (Ilse, Tomczak, and Welling 2018) as the aggregator for the vision branch, serving as a Baseline. Its performances by jointly training on all datasets (JointTrain, non-incremental) are taken as upper bound, and those by naively fine-tuning sequential datasets (FineTune) are taken as lower bound. In addition, we compare Regularization-based methods: EWC (Kirkpatrick et al. 2017) and LwF (Li and Hoiem 2017); Rehearsal-based methods: A-GEM (Chaudhry et al. 2018), ER (Chaudhry et al. 2019), DER++ (Buzzega et al. 2020), ER-ACE (Caccia et al. 2021) and ConSlide (Huang et al. 2023a); and Pre-trained VLM-based methods: AttriCLIP (Wang et al. 2023) and MI-Zero (Lu et al. 2023). To ensure a fair comparison, all compared methods are adapted to VLM. Specially, they use the same instance features and text encoder $\boldsymbol{E}_{\text{txt}}$ in CONCH.
@@
FineTune (Lower) 0.234 $\pm$ 0.008 0.262 $\pm$ 0.010 0.927 $\pm$ 0.023 -0.927 $\pm$ 0.023 0.803 $\pm$ 0.060
Regularization-based EWC (Kirkpatrick et al. 2017) 0.235 $\pm$ 0.011 0.264 $\pm$ 0.011 0.928 $\pm$ 0.020 -0.928 $\pm$ 0.020 0.833 $\pm$ 0.069
LwF (Li and Hoiem 2017) 0.236 $\pm$ 0.016 0.265 $\pm$ 0.012 0.908 $\pm$ 0.030 -0.908 $\pm$ 0.030 0.900 $\pm$ 0.041
Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.536 $\pm$ 0.047 0.591 $\pm$ 0.053 0.527 $\pm$ 0.067 -0.527 $\pm$ 0.067 0.872 $\pm$ 0.026
ER-ACE/30 (Caccia et al. 2021) 0.703 $\pm$ 0.049 0.777 $\pm$ 0.053 0.281 $\pm$ 0.062 -0.279 $\pm$ 0.063 0.889 $\pm$ 0.041
DER++/30 (Buzzega et al. 2020) 0.684 $\pm$ 0.055 0.755 $\pm$ 0.063 0.310 $\pm$ 0.069 -0.307 $\pm$ 0.072 0.910 $\pm$ 0.043
ER/30 (Chaudhry et al. 2019) 0.644 $\pm$ 0.028 0.711 $\pm$ 0.033 0.387 $\pm$ 0.050 -0.386 $\pm$ 0.049 0.901 $\pm$ 0.035
@@
Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.694 $\pm$ 0.058 0.766 $\pm$ 0.061 0.207 $\pm$ 0.063 -0.207 $\pm$ 0.063 0.861 $\pm$ 0.019
@@
(1) Our QPMIL-VL method is consistently superior to all other methods, with parameters only slightly larger than the smallest one. PQMIL-VL achieves an ACC of $0.890$$\pm$ $0.021$ with an Upper-bound Ratio as high as $0.982$$\pm$ $0.032$, winning SOTA performance metrics. Its Upper-bound Ratio is even $5.5\%$ higher than that of MI-Zero, the second-best method. Additionally, our model has a small learnable parameter size of $0.365$M, which is only slightly larger than the smallest size of $0.111$M in AttriCLIP. However, in terms of performance metrics, our method is far superior to the lowest-parameter method, i.e., AttriCLIP. Besides, in task-incremental experiments, our Masked ACC could reach up to $0.930$$\pm$ $0.018$, which is also significantly higher than any achieved by other compared methods.
@@
(1) Our QPMIL-VL method is consistently superior to all other methods in both class-incremental and task-incremental learning scenarios, maintaining a SOTA position. (2) Regularization-based methods are still ineffective in class-incremental learning scenario due to suffering from severe catastrophic forgetting. (3) Most rehearsal-based and pre-trained VLM-based methods often achieve good classification performance. Except for MI-Zero and our PQMIL-VL, the best ACC among these two types of methods reaches up to $0.703$$\pm$ $0.049$ with an Upper-bound Ratio of $0.777$$\pm$ $0.053$, achieved by ER-ACE.
Comparison between Forward-order and Reverse-order Training. Additionally, we observe that on reverse-order training, the performance of most methods is noticeably lower than on forward-order training. For example, our PQMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$. In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Table 1). One possible explanation to this decline is that in incremental learning, later datasets can affect earlier ones. ESCA, with only $150$ slides, is the first to be learned, while NSCLC, with $965$ slides, is the last. NSCLC is dominant in the total number of training samples, which could cause a greater influence on the earlier datasets, leading to more forgetting.
@@
I) Effect of Hyper-parameters As shown in Fig. 3, there are three important parameters in the prototype pool. The results indicate that inappropriate prototype pool capacity ((a): $N=7$, (b): $M<20$) may result in knowledge interference and too few prompts ((a): $N=1$) makes it difficult to describe the various instance prototype visual features present in WSIs.
@@
Figure 5: Class features visualization, only one class per dataset. Red $\star$ is the center, i.e., the average of features, for the instances of interest (orange points, determined by the cosine similarity between instance features and learned prototype features). Green $\circ$ and $\bullet$ denote the cosine similarities before and after feature enhancement, respectively.

### v2

@@ Abstract
Whole Slide Image (WSI) classification has very significant applications in clinical pathology, e.g., tumor identification and cancer diagnosis. Currently, most research attention is focused on Multiple Instance Learning (MIL) using static WSI datasets. One of the most obvious weaknesses of these methods is that they cannot efficiently preserve and utilize previously learned knowledge. With any new data arriving, classification models are required to be re-trained on both previous and current new data. To overcome this shortcoming and break through traditional vision modality, this paper proposes the first Vision-Language-based framework with Queryable Prototype Multiple Instance Learning (QPMIL-VL) specially designed for incremental WSI classification. This framework mainly consists of two information processing branches: one is for generating bag-level features by prototype-guided aggregation of instance features, while the other is for enhancing class features through a combination of class ensemble, tunable vector and class similarity loss. The experiments on four public WSI datasets demonstrate that our QPMIL-VL framework is effective for incremental WSI classification and often significantly outperforms other compared methods, achieving state-of-the-art (SOTA) performance. Our source code is publicly available at https://github.com/can-can-ya/QPMIL-VL.
@@ Introduction
Histopathology Whole Slide Image (WSI) is crucial for the diagnosis and treatment of tumor diseases (Song et al. 2023). To model gigapixel WSIs (e.g., $000\times 90,$ 000 × 90 , 000 pixels) for training clinical-grade models, many methods based on Multiple Instance Learning (MIL) have been extensively studied in the field of computational pathology (CPATH) (Campanella et al. 2019; Lu et al. 2021; Lin et al. 2023; Liu et al. 2024c). A shared characteristic of these methods is that they are specially designed for capturing a static data distribution (e.g., a given WSI dataset with two lung cancer subtypes), still following conventional machine learning paradigm. However, in the real-world, the distribution of WSI data could be dynamic due to the arrival of new datasets or the discovery of emerging cancer types (Van der Laak, Litjens, and Ciompi 2021; Derakhshani et al. 2022). As a result, a well-trained WSI classification model that is suitable for previous data often cannot adapt to new data. Therefore, it must be re-trained on the entire dataset (including both previous and new data), leading to a significant increase in training costs. This presents the necessities of new methods well-tailored to dynamic data distributions.
@@
Recent works, e.g., PLIP (Huang et al. 2023b), Prov-GigaPath (Xu et al. 2024) and CONCH (Lu et al. 2024), have deeply demonstrated the significant potential of Vision-Language Models (VLMs) in CPATH. Breaking through traditional vision modality, this kind of new models could effectively exploit pathology features from both vision and language modalities. All of these motivate us to leverage foundational pathology VLMs to study incremental learning for WSI classification.
@@
III) Extensive experiments are performed on four public WSI datasets. The results demonstrate the superiority of our QPMIL-VL over the existing methods.
@@
Prototype Pool. Specifically, the prototype pool consists of $M$ (key, prompt) prototype pairs $\{(\boldsymbol{k}_{i},\boldsymbol{P}_{i})\}_{i=1}^{M}$ initialized using a Normal Distribution, $\boldsymbol{k}_{i}\in\mathbb{R}^{D_{f}},\boldsymbol{P}_{i}\in\mathbb{R}^{L_{\boldsymbol{P}}\times D_{e}}$, where $D_{f}$ is the output feature dimension of encoder $\boldsymbol{E}_{\text{img}}$ and $\boldsymbol{E}_{\text{txt}}$, $D_{e}$ represents the embedding dimension of $\boldsymbol{E}_{\text{txt}}$ and $L_{\boldsymbol{P}}$ indicates the total number of learnable vectors in $\boldsymbol{P}_{i}$. In a prototype pair, the key works as a prototype identifier, used by WSIs for querying, while the prompt acts as a prototype descriptor, utilized to describe the specific visual feature of instance prototype within WSIs.
@@
where $Top\text{-}N^{min}$ denotes the operation of choosing the top-$N$ prototype keys with minimal $\boldsymbol{z}$-key distances. However, in prototype matching on different WSI datasets during incremental training by Eq. 1, we observe that almost all samples tend to choose the same set of $N$ prototypes from pool. This could usually lead to catastrophic forgetting in incremental learning. To effectively avoid this tendency, we define a penalty factor $p_{i}^{t_{c}}$ to optimize prototype choosing:
@@
FineTune (Lower) 0.308 $\pm$ 0.045 ∗∗∗ 0.336 $\pm$ 0.045 0.841 $\pm$ 0.054 -0.841 $\pm$ 0.054 0.844 $\pm$ 0.038 ∗∗∗
Regularization-based EWC (Kirkpatrick et al. 2017) 0.309 $\pm$ 0.051 ∗∗∗ 0.338 $\pm$ 0.055 0.836 $\pm$ 0.064 -0.836 $\pm$ 0.064 0.502 0.840 $\pm$ 0.035 ∗∗∗
LwF (Li and Hoiem 2017) 0.378 $\pm$ 0.081 ∗∗∗ 0.412 $\pm$ 0.083 0.731 $\pm$ 0.116 -0.731 $\pm$ 0.116 0.916 $\pm$ 0.031
Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.436 $\pm$ 0.058 ∗∗∗ 0.477 $\pm$ 0.070 0.670 $\pm$ 0.072 -0.670 $\pm$ 0.072 0.502 0.879 $\pm$ 0.045 ∗∗
ER-ACE/30 (Caccia et al. 2021) 0.666 $\pm$ 0.049 ∗∗∗ 0.732 $\pm$ 0.051 0.075 $\pm$ 0.057 -0.070 $\pm$ 0.059 0.917 $\pm$ 0.023
DER++/30 (Buzzega et al. 2020) 0.749 $\pm$ 0.055 ∗∗∗ 0.823 $\pm$ 0.059 0.219 $\pm$ 0.059 -0.214 $\pm$ 0.062 0.904 $\pm$ 0.029 ∗∗
ER/30 (Chaudhry et al. 2019) 0.790 $\pm$ 0.040 ∗∗∗ 0.870 $\pm$ 0.049 0.186 $\pm$ 0.044 -0.186 $\pm$ 0.044 0.894 $\pm$ 0.036 ∗∗
@@
ER/100 (Chaudhry et al. 2019) 0.827 $\pm$ 0.028 ∗∗∗ 0.910 $\pm$ 0.034 0.143 $\pm$ 0.034 -0.142 $\pm$ 0.032 0.502 0.918 $\pm$ 0.028
Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.616 $\pm$ 0.056 ∗∗∗ 0.677 $\pm$ 0.071 0.285 $\pm$ 0.059 -0.285 $\pm$ 0.059 0.111 0.844 $\pm$ 0.026 ∗∗∗
MI-Zero (Lu et al. 2023) 0.839 $\pm$ 0.034 ∗∗∗ 0.927 $\pm$ 0.044 - - - 0.909 $\pm$ 0.019 ∗∗
QPMIL-VL (ours) 0.890 $\pm$ 0.021 0.982 $\pm$ 0.032 0.027 $\pm$ 0.014 -0.027 $\pm$ 0.014 0.365 0.930 $\pm$ 0.018
Table 1: Main results in class-incremental learning scenario on forward-order training. The best performances are highlighted as bold. “Params” represents the number of learnable parameters. Gray numbers are directly cited from published paper. “/30” and “/100” mean buffer $30$ WSIs and buffer $100$ WSIs, respectively. We present the Masked ACC reflecting task-incremental learning scenario for reference. /∗∗/∗∗∗ denote there are significant differences (paired t-test p-value $<$ $0.05$/$0.01$/$0.001$) between the best performance and others.
@@
To optimize our QPMIL-VL model in training, based on the three losses above, total loss $\mathcal{L}_{T}$ is computed by
$\mathcal{L}_{T}=\mathcal{L}_{C}+\lambda\cdot\mathcal{L}_{M}+\beta\cdot\mathcal{L}_{S},$ (10)
where balance factors $\lambda$ and $\beta$ are set to constant $0.5$. Therefore, our optimization objective is to minimize $\mathcal{L}_{T}$.
@@
Following the setup of previous representative work (Van de Ven and Tolias 2019), we mainly evaluate QPMIL-VL in the class-incremental learning scenario and additionally report the results in the task-incremental learning scenario.
@@
Compared Methods. It is necessary to compare with the upper and lower bounds of IL methods. We employ the classic ABMIL (Ilse, Tomczak, and Welling 2018) as the aggregator for the vision branch, serving as a Baseline. Its performances by jointly training on all datasets (JointTrain, non-incremental) are taken as upper bound, and those by naively fine-tuning sequential datasets (FineTune) are taken as lower bound. In addition, we compare Regularization-based methods: EWC (Kirkpatrick et al. 2017) and LwF (Li and Hoiem 2017); Rehearsal-based methods: A-GEM (Chaudhry et al. 2018), ER (Chaudhry et al. 2019), DER++ (Buzzega et al. 2020), ER-ACE (Caccia et al. 2021) and ConSlide (Huang et al. 2023a); and Pre-trained VLM-based methods: AttriCLIP (Wang et al. 2023) and MI-Zero (Lu et al. 2023). To ensure a fair comparison, all methods are adapted to VLM. Specially, they use the same instance features and text encoder $\boldsymbol{E}_{\text{txt}}$ in CONCH.
@@
FineTune (Lower) 0.234 $\pm$ 0.008 ∗∗∗ 0.262 $\pm$ 0.010 0.927 $\pm$ 0.023 -0.927 $\pm$ 0.023 0.803 $\pm$ 0.060 ∗∗∗
Regularization-based EWC (Kirkpatrick et al. 2017) 0.235 $\pm$ 0.011 ∗∗∗ 0.264 $\pm$ 0.011 0.928 $\pm$ 0.020 -0.928 $\pm$ 0.020 0.833 $\pm$ 0.069 ∗∗∗
LwF (Li and Hoiem 2017) 0.236 $\pm$ 0.016 ∗∗∗ 0.265 $\pm$ 0.012 0.908 $\pm$ 0.030 -0.908 $\pm$ 0.030 0.900 $\pm$ 0.041
Rehearsal-based A-GEM/30 (Chaudhry et al. 2018) 0.536 $\pm$ 0.047 ∗∗∗ 0.591 $\pm$ 0.053 0.527 $\pm$ 0.067 -0.527 $\pm$ 0.067 0.872 $\pm$ 0.026 ∗∗∗
ER-ACE/30 (Caccia et al. 2021) 0.703 $\pm$ 0.049 ∗∗∗ 0.777 $\pm$ 0.053 0.281 $\pm$ 0.062 -0.279 $\pm$ 0.063 0.889 $\pm$ 0.041
DER++/30 (Buzzega et al. 2020) 0.684 $\pm$ 0.055 ∗∗∗ 0.755 $\pm$ 0.063 0.310 $\pm$ 0.069 -0.307 $\pm$ 0.072 0.910 $\pm$ 0.043
ER/30 (Chaudhry et al. 2019) 0.644 $\pm$ 0.028 ∗∗∗ 0.711 $\pm$ 0.033 0.387 $\pm$ 0.050 -0.386 $\pm$ 0.049 0.901 $\pm$ 0.035
@@
Pre-trained VLM-based AttriCLIP (Wang et al. 2023) 0.694 $\pm$ 0.058 ∗∗∗ 0.766 $\pm$ 0.061 0.207 $\pm$ 0.063 -0.207 $\pm$ 0.063 0.861 $\pm$ 0.019 ∗∗∗
@@
(1) Our QPMIL-VL method is consistently superior to all other methods, with parameters only slightly larger than the smallest one. QPMIL-VL achieves an ACC of $0.890$$\pm$ $0.021$ with an Upper-bound Ratio as high as $0.982$$\pm$ $0.032$, winning SOTA performance metrics. Its Upper-bound Ratio is even $5.5\%$ higher than that of MI-Zero, the second-best method. Additionally, our model has a small learnable parameter size of $0.365$M, which is only slightly larger than the smallest size of $0.111$M in AttriCLIP. However, in terms of performance metrics, our method is far superior to the lowest-parameter method, i.e., AttriCLIP. Besides, in task-incremental experiments, our Masked ACC could reach up to $0.930$$\pm$ $0.018$, which is also significantly higher than any achieved by other compared methods.
@@
(1) Our QPMIL-VL method is consistently superior to all other methods in both class-incremental and task-incremental learning scenarios, maintaining a SOTA position. (2) Regularization-based methods are still ineffective in class-incremental learning scenario due to suffering from severe catastrophic forgetting. (3) Most rehearsal-based and pre-trained VLM-based methods often achieve good classification performance. Except for MI-Zero and our QPMIL-VL, the best ACC among these two types of methods reaches up to $0.703$$\pm$ $0.049$ with an Upper-bound Ratio of $0.777$$\pm$ $0.053$, achieved by ER-ACE.
Comparison between Forward-order and Reverse-order Training. Additionally, we observe that on reverse-order training, the performance of most methods is noticeably lower than on forward-order training. For example, our QPMIL-VL decreases to an ACC of $0.859$$\pm$ $0.032$. In contrast, its ACC is $0.890$$\pm$ $0.021$ on forward-order training (as shown in Table 1). One possible explanation to this decline is that in incremental learning, later datasets can affect earlier ones. ESCA, with only $150$ slides, is the first to be learned, while NSCLC, with $965$ slides, is the last. NSCLC is dominant in the total number of training samples, which could cause a greater influence on the earlier datasets, leading to more forgetting.
@@
I) Effect of Hyper-parameters As shown in Fig. 3, there are three important parameters in the prototype pool. The results indicate that inappropriate prototype pool capacity ((a): $N=7$, (b): $M<20$) may result in knowledge interference and too few prompts ((a): $N=1$) make it difficult to describe the various instance prototype visual features present in WSIs.
@@
Figure 5: Class features visualization, only one class per dataset. Red $\star$ is the center, i.e., the average of features, for the instances of interest (orange points, determined by the cosine similarity between instance features and learned prototype features). Blue points are non-interest instance features. Green $\circ$ and $\bullet$ denote the cosine similarities before and after feature enhancement, respectively.